OpenStudy (anonymous):

can some one help me .. -6 - 8c = 3c+16

OpenStudy (anonymous):

\[-6-8c=3c+16\]

OpenStudy (jennychan12):

-6 - 8c = 3c+16 -11c = 22 c = 2 You just have to brush up on your algebra if you have trouble with this.

OpenStudy (anonymous):

thank you very much